OpenStudy (anonymous):

A baseball minor league's ticket price decreased $2.50 for each of 2 years, but then increased $4.25 in each of the next 3 years. What was the overall increase of the ticket price? A. $7.75 B. $12.75 C. $3.50

OpenStudy (anonymous):

(2 years) x (-2.50) + (3 years) (4.25) = -5.00 + 12.75 = 7.75

OpenStudy (anonymous):

its A i think im not sure on this one